### v1

# Deep Learning Reduced Order Modelling on Parametric and Data driven domains

Abstract
Partial differential equations (PDEs) are extensively utilized for modeling various physical phenomena. These equations often depend on certain parameters, necessitating either the identification of optimal parameters or solving the equation across multiple parameters to understand how a structure might react under different conditions. Performing an exhaustive search over the parameter space requires solving the PDE multiple times, which is generally impractical. To address this challenge, Reduced Order Models (ROMs) are constructed from a snapshot dataset comprising parameter-solution pairs. ROMs facilitate the rapid solving of PDEs for new parameters.
Recently, Deep Learning ROMs (DL-ROMs) have been introduced as a new method to obtain ROM. Additionally, the PDE of interest may depend on the domain, which can be characterized by parameters or measurements and may evolve with the system, requiring parametrization for ROM construction. In this paper, we develop a Deep-ROM capable of extracting and efficiently utilizing domain parametrization. Unlike traditional domain parametrization methods, our approach does not require user-defined control points and can effectively handle domains with varying numbers of components. Moreover, our model can derive meaningful parametrization even when a domain mesh is unavailable, a common scenario in biomedical applications. Our work leverages Deep Neural Networks to effectively reduce the dimensionality of the PDE and the domain characteristic function.
Keywords Reduced Order Modeling  $\cdot$ Deep Learning  $\cdot$ Partial Differential Equations
1 Introduction
Applications such as design optimization and real-world simulation often require solutions to partial differential equations across various parameters [1], [2], [3], [4]. Typically, with given parameters $\bm{\lambda}$ the objective is to find a function $\bm{u}$ such that:
$\mathcal{L}(\bm{u};\bm{\lambda})=f(\cdot;\bm{\lambda}),\text{ on $\Omega(\bm{\lambda})$}$ (1)
$\mathcal{B}(\bm{u};\bm{\lambda})=g(\cdot;\bm{\lambda}),\text{ on $\partial\Omega(\bm{\lambda})$},$ (2)
where $\mathcal{L}$ and $\mathcal{B}$ are differential operators, $f$ and $g$ are functions and $\Omega$ is the domain, and all of these can depend on the parameter vector $\bm{\lambda}$. This task can be formulated as finding a mapping from parameters to their corresponding solutions. A naive approach would be to just solve PDE for each parameter value we want to examine. This can often be prohibitive due to time and memory constraints. To mitigate this problem, a Reduced Order Model (ROM) is constructed, which relies on the assumption that all solutions for a specified parametric PDE belong to a lower-dimensional space. Classically, the problem of constructing ROM then involves determining the reduced basis [5]. In order to construct the ROM, one needs to obtain a dataset of high-fidelity parameter-solution pairs. This is usually done by employing the Galerkin finite element method (FEM) [6] which approximates solutions within a finite-dimensional space and uses a weak formulation. To solve a single instance of a PDE, the solution is approximated in a finite-dimensional function space. The problem of solving the PDE is then reformulated as the problem of finding the coefficients of the basis functions in that space. This results in a large system of equations and the discretized solution of PDE can be written as a large vector. To solve for many parameter values, one constructs and uses ROM. Its development consists of two phases: offline phase in which the ROM is constructed, and online phase, where the solution for the new parameter is evaluated. To construct ROM, one collects the dataset $(\bm{\lambda},\bm{u}_{i}),i=1,\ldots,N$, where $\bm{\lambda}$ is the i-th vector of parameters and $\bm{u}_{i}$ is the full-order model (FOM) solution corresponding to these parameters. The dataset comes from measurements or high-fidelity numerical solvers. Usually, using this data, linear reduction techniques are used to project solutions and equations into lower-dimensional space to obtain the ROM. To get a solution for the new parameter, it is then only necessary to solve in lower-dimensional space, which is faster than solving the original PDE. However, evaluation still requires solving a smaller system of equations. Furthermore, classical ROMs depend on the quality of the linear projection. This may result in a high number of equations for a reduced model. Because of that, few nonlinear methods were developed [7], [8], [9].
Regarding classical Reduced Order Models (ROMs), constructing ROMs that are effective across different domains presents several challenges. Firstly, to identify the projection onto the appropriate lower-dimensional space, all components of all solutions must have matching locations on the computational domain. This requirement is often unfeasible because irregular domains are discretized differently, making it impossible to ensure that all solutions have the same dimensionality. A common solution to this issue is to create a mapping between a fixed reference domain and the domain of interest, transforming the equation to apply to the fixed reference domain [10]. However, this approach is highly problem-dependent and not universally applicable. Additionally, domains may be derived from measurements, have varying numbers of subdomains or holes, or lack an available mesh. In such cases, it is unclear how to parametrize these domains and construct an appropriate mapping to a reference domain.
In this work, we address these challenges by employing convolutional autoencoder architecture on characteristic functions represented as a bitmap image to extract domain parametrization. This approach works in cases where the mesh is not available, when domains come from measurements, or when we have varying numbers of components or holes. The extracted parametrization is then incorporated into a Deep-ROM architecture. We demonstrate the effectiveness of our approach across various 2D problems.
2 Related work
Recently, there has been an increasing interest in using Deep Learning for Reduced Order Modelling. Since neural networks are universal approximators [11], [12], [13], it follows that they can also approximate the solution map. In this section, various approaches that use deep learning for the construction of the ROM, as well as some classical approaches, are described. Throughout this section, it is assumed that the training set of parameter-solution pairs $(\bm{\lambda}_{i},\bm{u}_{i}),i=1,2,\ldots N$ is obtained via measurements or a numerical method such as the Galerkin finite element method [14].
2.1 Reduced Order Model via Proper orthogonal decomposition
Proper orthogonal decomposition (POD) [15] is a linear reduction technique. The basic assumption is that even though discretized solutions belong to a high-dimensional vector space of dimension $m$, solutions can be well approximated by a lower-dimensional vector space of dimension $k\ll n$. In this setting, all solutions can be vectorized and assembled into a matrix $\mathbf{U}\in\mathbf{M^{mn}}$. To obtain the basis of such a lower-dimensional vector space, SVD decomposition [16] is calculated, which means that orthogonal matrices $\mathbf{\hat{U}}\in\mathbf{M^{m}},\mathbf{\hat{V}}\in\mathbf{M^{n}}$ orthogonal matrices and diagonal $\Sigma\in\mathbf{M^{mn}}$ with non-negative decreasing numbers on its diagonal $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{m}$ are found such that
$\mathbf{U}=\mathbf{\hat{U}\Sigma\hat{V}}.$
The matrix of orthogonal basis vector for lower-dimensional vector space $V$ is then defined by extracting the first $k$ columns of matrix $\mathbf{\hat{U}}$.
We can then assume that any solution of this parametric problem can be written as
$\mathbf{u}=\mathbf{Vx},$
where $\bm{x}$ is $k$-dimensional vector. Now, instead of finding $\bm{u}$, we have to find $\bm{x}$. Then, when $Vx$ is substituted instead of $\bm{u}$ in the Galerkin FOM formulation, a new system with fewer unknowns is obtained. In Linear PDE, this results in decreasing the number of equations. In non-linear cases, this cannot be trivially done, so techniques such as the Discrete Empirical Interpolation Method (DEIM) [17] are developed. Further difficulties in employing this ROM arise when the domain depends on parameters because all discretized solutions need to have the same dimensionality and each coordinates of all solutions have to correspond to the same points in the domain. To obtain ROM on parametric domains $\Omega(\bm{\lambda})$, equations are mapped on reference domain $\hat{\Omega}$. Mapping $\mathcal{M}(\cdot;\bm{\lambda})$ maps the reference domain into $\Omega(\bm{\lambda})$, so the correspondence of points across different domains is established. To construct such a mapping, one can use free-form deformation (FFD) [18], [19], [20], radial basis function interpolation (RBF) [21], [22], inverse distance weighting interpolation (IDW) [23], [24], [25], or surface registration using currents [26]. Furthermore, domains should be parametrized as well. To get parametrization, a set of control points $\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{c}$ around the domain is selected, and displacements of these points from their corresponding positions on the reference domain are calculated. Control points on each domain have to match the same points in the reference domain through appropriate mappings to get meaningful parameters. Control points have to be chosen carefully so that they capture the shape characteristics well. This imposes difficulties when one wants to obtain one ROM for situations where there are a varying number of components or holes.
2.2 Physical Informed Neural Networks for Reduced Order Model
In [27], the authors use a different approach to construct the ROM. They construct the ROM via POD. The goal of the model is to predict coefficients in the reduced basis from parameter values using Physics-Informed Neural Networks (PINNs) [28]. The loss function is the sum of the mean-squared error between predicted and correct coefficients and the mean-squared error of the equation residual of the reduced system. If not enough high-fidelity solutions are available, the model can be trained to minimize the equation residual loss only.
2.3 Deep ROM (DL-ROM)
N. Franco et al. introduced Deep ROM (DL-ROM) in [29]. The idea is to obtain ROM by exploiting autoencoder architecture [30]. The model consists of three neural networks: an encoder $\Psi_{E}^{S}$, a decoder $\Psi_{D}^{S}$, and a multi-layer perceptron (MLP) [31] $\Phi_{S}$. To obtain ROM, the encoder-decoder pair is trained to minimize reconstruction error:
$\frac{1}{N}\sum_{i=1}^{n}\|\bm{u}_{i}-\Psi_{D}^{S}(\Psi_{E}^{S}(\bm{u}_{i}))\|^{2}.$
The encoder for standard DL-ROM $\Psi_{E}^{S}(\bm{u}_{i})$ maps the solution to a lower-dimensional space, called encoding, and the decoder lifts encoding to the original solution size. The autoencoder is then trained to minimize reconstruction error. Finally, $\Phi_{S}$ is trained to minimize the mean-squared error between its output and encoder encoding:
$\frac{1}{N}\sum_{i=1}^{N}\|\Phi_{S}(\bm{\lambda}_{i})-\Psi_{E}^{S}(\bm{u}_{i})\|^{2}.$
Solution $\bm{u}$ for a new value of parameter $\bm{\lambda}$ is then:
$\bm{u}=\Psi_{D}^{S}(\Phi_{S}(\bm{\lambda})).$ (3)
The whole architecture is shown on Figure 1
Figure 1: Basic DL-ROM architecture: encoder $\Psi_{E}^{S}$ and decoder $\Psi_{D}^{S}$ are trained to minimize reconstruction error between full order model solution and its reconstruction, After training, encodings of solutions are calculated with encoder $\Psi_{E}^{S}$ and Multilayer perceptron (MLP) $\Phi_{S}$ is trained to map parameters to encoding. To obtain a solution for new parameters $\bm{\lambda}$, parameters are passed through $\Phi_{S}$ and then through the decoder $\Psi_{D}^{S}$.
Notice that getting solutions from this kind of reduced-order model doesn’t require solving any additional equations and can be efficiently done for multiple parameters at the same time, suing GPU parallel computing. In this work, Convolutional Neural Networks (CNN) [32] were used for the encoder and decoder and all domains involved were rectangular.
2.4 Deep Learning Proper Orthogonal Decomposition ROM
Deep Learning Proper Orthogonal Decomposition ROM (POD-DL-ROM) was introduced in [33]. The procedure goes as follows. Solutions $\bm{u}_{i},i=1,\ldots N$ are reshaped into vectors of length $N_{h}$ and assembled into a matrix $\mathbf{A}$ with $N_{h}$ rows and $N$ columns. Columns of matrix $\mathbf{A}$ are $\bm{a}_{i},i=1,2,\ldots N$. The essence of the POD method is to get orthonormal vectors $\bm{\Phi_{i}},i=1,2,\ldots k$, $k\ll N$ vectors that maximize:
$\frac{1}{N}\sum_{i=1,\ldots N,j=1,\ldots k}|\langle\bm{a}_{i},\bm{\Phi}_{j}\ \rangle|^{2}$
So, the following approximation hold:
$\bm{u}\approx\mathbf{V}_{k}\bm{u}_{k}$
Matrix $\mathbf{V}_{k}$ is usually found using SVD decomposition or randomized SVD (rSVD) [34]. So, the solution $u$ can be approximated with POD basis and represented as $k$-dimensional vector $[\langle\bm{u},\bm{\Phi}_{1}\rangle,\ldots\langle\bm{u},\bm{\Phi}_{k}\rangle]$. Then, similar to the black-box DL-ROM, the model consists of three networks: encoder $\Psi_{E}$, decoder $\Psi_{D}$, and parameter-encoding network $\Phi$. The difference is that the encoder and decoder are trained on POD coefficients, thus reducing the dimension of the problem. For this model to work on multiple domains, POD coefficients at the reference domain and mapping between domain and reference domain are needed.
2.5 Graph convolutional autoencoder approach for DL-ROMs
Following DL-ROM framework, in [35] Graph convolutional autoencoders (GNNs) [36], [37], [38], [39] are used for dimension reduction. Using the proposed autoencoder technique, one can obtain ROM for non-rectangle domains using the GNN autoencoder. In the proposed approach, graph pooling and a fully-connected layer work only when all domains have the same number of nodes and domain parametrization is known as well. In applications, especially in biomedicine, both constraints may not hold.
3 Our method
In this work, we develop an efficient Reduced Order Model (ROM) that operates effectively across different domains, based on a Deep Learning Reduced Order Model (DL-ROM) approach. We refer to our method as Deep Learning domain aware ROM (DL-DA ROM). To achieve this, meaningful parametrization of domains is essential. Domains can be described parametrically, may evolve as part of the system (e.g., moving domains), or can be obtained from measurements.
Since the ROM for given parameters has to predict the domain and values of solutions on that domain, we separate those two tasks. In this work, we assume that the solution is at least continuous and can be interpolated onto rectangular mesh without significant loss of accuracy, see Figure 2. By interpolating in this manner, Convolutional neural networks (CNNs) can be utilized to construct an autoencoder for solutions. This approach is particularly beneficial in scenarios where the domain is derived from measurements and meshing is prohibitively expensive.
Figure 2: Sample interpolated onto equidistant rectangular grid (on the left). Sample (on the right).
When domains are parametrically described, it is not necessary to assume that each domain has the same number of parameters; they may have varying numbers of subdomains or holes. Our approach assumes that each domain characteristic function can be interpolated onto a rectangular grid, see Figure 3.
Figure 3: Mesh representation of the domain (left), an image of domain characteristic function (right).
The domain of the problem can be derived either from measurements or parameters. In our proposed method, its characteristic function is described as a bitmap image, with pixels assigned a value of one or zero depending on whether the pixel is inside or outside the domain. A convolutional autoencoder is trained using binary cross-entropy loss [40] to classify each pixel as part of the domain or not which is given by
$l(x,y)=\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}-y_{ij}\log(x_{ij})-(1-y_{ij})\log(1-x_{ij}),$
where $\bm{y}$ is target bitmap image with $n$ rows and $m$ columns and $\bm{x}$ is network output image, whose pixels values $x_{ij}$ denote probability that pixel $x_{ij}$ is equal to one. The encoded domains are then used as the parametrization of the domain. The domain autoencoder is inspired by semantic segmentation encoder-decoder models in computer vision [41], [42], with two classes indicating whether a pixel is inside or outside the domain. This approach can be extended to problems with multiple domains, such as fluid-structure interaction problems, providing simultaneous geometric parametrization for the entire system. If domains can be generated synthetically, it is possible to use more samples for the domain autoencoder than for the solution autoencoder.
Domain autoencoder is shown in Figure 4.
Figure 4: Domain Autoencoder used for domain parametrization. Input to the convolutional neural network is a characteristic function represented as a bitmap image. The output of the autoencoder is a rectangular grid with the probability that a given pixel belongs to the domain. The encoder output is used as domain parametrization.
If the domain is either known from measurements or is parametrized, the encoding of the domain can be used as additional parameters in the network $\Phi$. The model is shown in Figure 5 and all important terms in this work are given in table 1.
[sca]
Figure 5: Deep ROM for parametric domains: In addition to vanilla Deep ROM, we introduce domain autoencoder, which extracts parameters for the domain. These encodings are then used as additional parameters in Deep ROM.
label meaning $\Omega(\bm{\lambda})$ domain of the problem, in our work, is dependent on parameters $\Psi_{E}^{S}$ solution encoder $\Psi_{D}^{S}$ solution decoder $\Phi_{S}$ neural network that maps parameters into solution encoding $\Psi_{E}^{D}$ domain encoder $\Psi_{D}^{D}$ domain decoder
Table 1: Notation used in this work.
Constructing the ROM, also known as the offline phase, involves gathering the data, either with numerical simulation, measurements, or possibly a combination of both. Domain and solution autoencoders are trained to obtain meaningful low-dimensional encodings. Before training the solution autoencoder training solution data is standardized by applying the following transformation:
$\bm{u}_{i}=\frac{\bm{u}_{i}-\mu_{s}}{\sigma_{s}},1,\ldots N$
where $N$ is a number of samples in the training dataset, $\mu_{s}$ and $\sigma_{s}$ are mean and standard deviation of the training dataset After training the solution and domain autoencoders, domain encoding, and other equation parameters are concatenated. Training parameters are standardized by applying the following transformation:
$\lambda_{ij}=\frac{\lambda{ij}-\mu_{pj}}{\sigma_{pj}},i=1,\ldots N,j=1\ldots k$
where $k$ is dimension of parameter vector, $N$ number of samples and $\mu_{pj},\sigma_{pj}$ are mean and standard deviation of each parameter separately. Afterward, the network $\Phi_{S}$ is trained to predict solution encodings from parameters.
The procedure is summarized in Algorithm 1. After all data collection and training are completed, solutions for unseen parameters are calculated in the online phase, as detailed in Algorithm 2. For new values of equation parameters and domain, this involves passing the domain through the domain encoder $\Psi_{E}^{D}$ concatenating the resulting encoding with the equation parameters, passing them through the MLP $\Phi_{S}$ and finally through the solution decoder $\Psi_{D}^{S}$ to obtain prediction $\hat{u}$. The final ROM solution is then given by:
$\bm{u}=\sigma_{s}\bm{u}+\mu_{s}$
Algorithm 1 Offline phase
Parameter-solution-characteristic function triplets $(\bm{\lambda}_{i},\bm{u}_{i},\bm{c}_{i}),i=1,\ldots N$, hyperparameters for training of networks $\Psi_{E}^{S},\Psi_{D}^{S}$, domain autoencoder $\Psi_{E}^{D},\Psi_{D}^{D}$, and parameter-encoding network $\Phi_{S}$
Trained networks
Standardize solution data
Train solution encoder $\Psi_{E}^{S}$ and decoder $\Psi_{D}^{S}$ to minimize reconstruction loss $\frac{1}{N}\sum_{i=1}^{N}\|\bm{u}_{i}-\Psi_{D}^{S}(\Psi_{E}^{S}(\bm{u}_{i}))\|^{2}$
Evaluate trained $\Psi_{E}^{S}$ on $\bm{u}_{i},i=1,\ldots N$ and get encodings $\bm{e}_{i}=\Psi_{E}^{S}(\bm{u}_{i}),i=1,\ldots N$
Train domain encoder $\Psi_{E}^{D}$ and decoder $\Psi_{D}^{D}$ to minimize binary cross-entropy loss
Evaluate trained $\Psi_{E}^{D}$ and get domain parameters $\bm{d}_{i}=\Psi_{E}^{D}(\bm{c}_{i}),i=1,\ldots N$
Concatenate domain and equation parameters $\bm{d}_{i}$ and $\bm{\lambda}_{i}$ into new parameters $\bm{f}_{i},i=1,\ldots N$
Standardize new parameters $\bm{f}_{i},i=1,\ldots N$
Train $\Phi_{S}$ to minimize reconstruction loss $\frac{1}{N}\sum_{i=1}^{N}\|\Phi_{S}(\bm{f}_{i})-\bm{e}_{i}\|^{2}$
Save all trained networks
Save mean $\mu_{s}$ and standard deviation $\sigma_{s}$ of training dataset and mean $\mu_{p}$ and standard deviation $\sigma_{p}$ of training feature dataset
Algorithm 2 Online phase
: Neural networks $\Psi_{D}^{S},\Psi_{E}^{D}$, and $\Phi_{S}$, values of parameters and/or domains for which we want to evaluate solution $(\bm{\lambda}_{i},\bm{c}_{i}),i=1,\ldots M$, mean and standard deviation of training parameters $\bm{f}_{i}$ $\mu_{p},\sigma_{p}$, mean and standard deviation of solutions $\mu_{s},\sigma_{s}$
Solutions for given parameters and domains $\bm{u}_{i},i=1,\ldots M$
Compute domain parameters $\bm{d}_{i}=\Psi_{E}^{D}(\bm{c}_{i})$
Concatenate $f_{i}=(\bm{\lambda}_{i},\bm{d}_{i}),i=1,\ldots M$
Standardize $\bm{f}_{i}$ based on training mean and standard deviation
Calculate ROM prediction $\bm{\hat{u}}=\Psi_{D}^{S}(\Phi_{S}(\bm{f}_{i}))$
Return solution $\sigma_{s}\bm{\hat{u}}+\mu_{s}$
Experiments and Results
Through this section, we will use relative error to evaluate our approach. The relative error $\epsilon$ for sample $(\bm{\lambda},\bm{u}_{h},\bm{d}_{h})$ is:
$\epsilon(\bm{u}_{h},\hat{\bm{u}_{h}})=\frac{\sqrt{\sum_{i,j}(u^{i,j}_{h}d^{i,j}_{h}-\hat{u}^{i,j}_{h}d^{i,j}_{h})^{2}}}{\sqrt{\sum_{i,j}(u^{i,j}_{h})^{2}d^{i,j}_{h}}},$
where $\bm{\lambda}$ is the parameter vector, $\bm{u}_{h}$ solution discretized on a mesh with $n\times m$ points, $\bm{\hat{u_{h}}}$ is DL-ROM prediction and $\bm{d}_{h}$ is an approximation of characteristic functions of the domains on that same mesh.
Relative error at i-th row and j-th column is:
$\epsilon_{i,j}(u_{k},\hat{u}_{h})=\frac{u_{h}^{i,j}d_{h}^{i,j}-\hat{u}^{i,j}_{h}d^{i,j}_{h}}{\sqrt{\sum_{k,l}(u^{k,l}_{h})^{2}d^{k,l}_{h}}},$
where $u^{i,j}_{h},\hat{u}^{i,j}_{h},d^{i,j}_{h}$ are discretized solutions, model approximation of the solution, and characteristic function value in i-th row and j-th column.
We use FEM implementation in FreeFEM [43] to obtain solutions and Pytorch [44] for defining and training neural networks. All solutions are projected on a rectangular equidistant mesh, in order to use CNN. In all examples, Leaky ReLU is used as an activation function after convolution. We use batch normalization [45] after each layer, except for the last layers of the encoder and decoder. In all examples, the number of channels per layer in the decoder mirrors the encoder architecture. Upsampling operations are used only in some layers, such that the resolution of outputs in the decoder layers mirrors the resolution of outputs in the encoder layers. For all networks, Adam optimizer [46] with One Cycle Policy scheduling [47] is used that adjusts the learning rate at each training step, with a maximum of 1e-3 for the encoder-decoder network.
All training, test, and validation sets used have $1000$ samples. The batch size for the autoencoder is $50$. Autoencoder network hyperparameters were chosen by trial and error with respect to autoencoder reconstruction, where we searched on latent dimensions $15,20,30$ and the number of channels for each encoder layer (the decoder mirrors the encoder). Tried hyperparameters are shown in Table 2. Autoencoders with more channels per layer had better training and test accuracy, so we picked the best two for further analysis.
Latent dimension $[15,20,30]$
Nbr. of kernels in encoder $[4,4,8,8,16,16,16]$, $[4,8,16,32,32,64,64]$ , $[8,16,16,32,64,64,64]$
Learning rate 1e-3
Batch size 50
Kernel size 5
Table 2: Autoencoder grid search hyperparameters. We trained 9 models with varying channel sizes and latent dimension sizes. The number of channels in decoder layers mirrors the encoder number of channels per layer. We have chosen two autoencoders with the smallest reconstruction errors on the test set, both with $8,16,16,32,64,64,64,64$ channels in encoders and dimensions $20$ and $30$ for further analysis.
We have not encountered any overfitting problems by increasing autoencoder capacity. However, neural network $\Phi_{S}$ easily overfits. Since $\Phi_{S}$ is usually small compared to autoencoder and trains fast, we employed grid search to choose the best hyperparameters for $\Phi_{S}$. Hyperparameters are shown in Table 3. The grid search was done for two ROM representations, obtained from solution autoencoders with $8,16,16,32,64,64,64$ channels per layer and latent dimensions equal to $15$ and $30$, respectively. Domain features are obtained from autoencoders trained on $50000$ domains with $32,32,64,64,64,64,64,64$ channels per layer in each layer and strides $1,1,2,2,2,1,1,1$ per layer respectively. For each model, the One Cycle Policy scheduler with Adam optimizer was used. We have trained two different domain autoencoders, one with latent dimension $15$ and one with latent dimension $30$. We choose the best hyperparameters with respect to the mean-squared error between the prediction of $\Phi_{S}$ and ROM representation.
Hyperparameter Values
Training set size 1000
Number of epochs 1500
Number of layers 1,2,3,4
Neurons per layer 64, 128, 256, 512, 1204, 2048
Dropout rate 0, 0.025, 0.05, 0.1, 0.2
Maximum learning rate 1e-5, 3.16e-5, 1e-4, 3.16e-4, 1e-3, 3.16e-3, 1e-2
Batch size $8,16,32,64,128$
Table 3: Hyperparameters grid search for $\Phi$ is performed, which results in total $4200$ models.
3.1 Advection-diffusion on different domains
Advection-diffusion equation on parametric domains:
$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ without elipse}\\ u=1&\text{ on edge of ellipse}\\ u=0&\text{for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (4)
Parameters are:
  • transport angle $\phi\in[0,2\pi]$ which determines vector $\mathbf{c}=(cos(\phi),sin(\phi))$
  • center of ellipse $(x_{0},y_{0})\in[0.1,0.9]^{2}$
  • angle of major axis of ellipse and x-axis $\alpha\in[0,\pi]$
  • $\beta\in[1,10]$
  • half-axes for ellipse are then taken randomly from interval depending on the center and angle of the ellipse
In total, there are $7$ parameters.
In this example, we know exact geometric parametrization, so we can obtain DL-ROM performance with calculated and exact parametrization and directly compare those two. For each domain parametrization, we have trained a fully-connected neural network $\Phi_{S}$ with hyperparameters determined by grid search.
Comparison of vanilla DL-ROM and our approach is shown in Table 4. We compare DL-ROM performance when exact ellipse parameters are used with domain parametrization calculated with the autoencoder. In all cases, models with calculated ellipse parameters perform better on the test set than models with exact ellipse parameters. In fact, we report that more than $1300$ models out of $4200$ trained in grid search, performed better on parameters of domain obtained autoencoder than the best model trained on exact domain parameters, for each combination. In this case, it can be seen that vanilla DL-ROM is more sensitive to ROM dimension, than our approach. Results are shown on Figure 6.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact ellipse parameters only 0.0554552 0.0554552 0.0605261 0.0605261 exact and calculated elipse parameters 0.0454250 0.0425800 0.0458577 0.0438216 calculated elipse parameters 0.0465742 0.0452926 0.0451946 0.0449190
Table 4: Comparison of DL-ROM and DL-DA ROM performance: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension. We differentiate between cases where geometric parameters are exactly known (1st row), when we add domain-encoded features to exact geometric features (2nd row), and when we use only geometric parameters obtained by autoencoder (3rd row). We can see that using autoencoder domain parametrization is beneficial for model performance. We report on the best model for each ROM encoding dimension and domain encoding dimension. Models trained on calculated domain parameters perform better than models trained using only exact domain parameters.
Figure 6: Top row: Exact solution, left, network prediction for a model trained on exact domain parameters (center), network prediction for a model trained on approximated domain parameters. Bottom row: relative error for a model trained on exact domain parameters (left), relative error for a model trained on approximated domain parameters (middle), relative error for a model trained on both exact and calculated domain parameters (right). Results are shown for ellipse with center $(0.76,0.38)$, half-axes $a=0.14,b=0.11$ and angle $\alpha=1.51$, $\beta=6.92$, $\mu=0.29$. All models were trained on $1000$ samples.
3.2 Domains with varying number of parameters
We experiment again on the advection-diffusion equation on parametric domains. In this case, domains can depend on a different number of parameters. Such a domain may be difficult to parametrize by classical techniques. We are investigating domains that are squares with $1$ to $4$ circular holes. Equations are:
$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ whitout circles}\\ u=2&\text{ on edge of circles}\\ u=0&\text{ for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (5)
Parameters are:
  • transport angle $\phi\in[0,2\pi]$ which determines vector $\mathbf{c}=(cos(\phi),sin(\phi))$
  • $m\in[1,4]$ number of circular holes
  • centers of holes $(x_{i},y_{i})$, $i=1,\ldots m$.
  • circles radii $r_{i}$, $i=1,\ldots m$.
  • $\beta\in[1,10]$
To generate domains, first, we randomly choose the number of circles $m$. Then we randomly choose the centers of circles so the distance between the centers and the square boundary is at least $0.1$. Finally, radii are chosen randomly so that the distance between all circles is at least $0.1$ and $r\in[0.05,0.2]$. This procedure ensures that circles do not overlap and that all circles are of reasonable size. We have chosen the same encoder-decoder architectures as in the previous example and after training solution autoencoders, a grid search was performed to obtain $\Phi_{S}$ hyperparameters. Results are shown in Table 5 and Figure 7. A larger ROM dimension mostly does not affect DL-ROM performance, while a smaller domain parametrization encoding dimension is beneficial in this case.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 calculated domain parameters 0.0239904 0.0264217 0.0209590 0.0267886
Table 5: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension. In this case, domains have varying numbers of parameters. We use only geometric parameters obtained by the autoencoder. We report on the best model for each ROM encoding dimension and domain encoding dimension.
Figure 7: Exact solutions (left), model prediction (middle), relative error (right) for model trained on $1000$ random samples for parameters $\beta=0.977$, $\mu=0.842$, with centers $(0.23,0.73),(0.57,0.22),(0.25,0.37),(0.79,0.79)$ and radii $0.11,0.10,0.14,0.10$ respectively.
Furthermore, we test how sensitive DL-DA ROM DL-ROM is on small perturbations of the circle boundary. The sensitivity of this model is especially important if we train on synthetically generated data and test only on possibly imprecisely measured domains and parameters. Instead of testing the trained model on a domain with circular holes, we test the model on predicting the correct solution when the domain is a unit square without ellipses. The idea is that these ellipses model slightly deformed circles. The correct solution is again generated using the finite element method. We tested on solutions with parameters $\beta=1.19617,\theta=1.71657$ with $4$ holes, with centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively. A comparison between FEM and our DL-ROM is shown in Figure 8. We tested how sensitive our model is to perturbations of the boundary. In these cases, instead of circular holes, we have elliptic holes. The relative error for increasing deformation is given in Table 6.
Ratio between small and large half-axis two half-axes in % $\epsilon$
1 0.0101
0.95 0.0120
0.9 0.0142
0.8 0.0166
Table 6: Relative errors for the same $\beta$, $\theta$ and ellipse centers as in 8 ($\beta=1.19617$, $\theta=1.71657$). Ellipses have centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$, radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively and varying rations between small and large half-axis. We can see an increase in relative error with higher deformation of the domain boundary, but the increase is not rapid.
Figure 8: Exact solutions (left) not multiplied by characteristic function, model prediction (middle) not multiplied by characteristic function, relative error (right) multiplied by characteristic function for slightly deformed circles with parameters $\beta=1.19617$, $\theta=1.71657$. Ellipses have centers, $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively. Small half-axes are $20\%$ smaller than large half-axes for each ellipse. Even though the model was not trained on elliptic holes, we can see a small relative error further from the boundary. The error is the largest at the boundary since the solution autoencoder did not capture ellipses in the original training dataset. This demonstrates that the model can work with small perturbations of the boundary.
Conclusion
In this work, we present the extension of DL-ROM to various 2D domains. This approach can be extended to 3D problems. Our algorithm can be used in problems where domains come from measurements. However, this comes at the expense of the main disadvantage of this approach, which is the first step of interpolating into a rectangular domain. In some problems, such as 3D thin pipes with torsion, rectangular data can be much larger than the actual number of nodes required to describe the pipe, which can be a computational bottleneck. In future work, we plan to mitigate this issue as well. We demonstrated that meaningful geometrical parametrization can be obtained using neural networks, which may be useful in other reduced-order model techniques as well.
Acknowledgments
This research was performed using the advanced computing service provided by the University of Zagreb University Computing Centre - SRCE. This research was supported by the Croatian Science Foundation under the project number IP-2022-10-2962 (I.M., B.M. and D.V.).

### v2

@@
-Deep Learning Reduced Order Modelling on Parametric and Data driven domains
+Reduced Order Modeling of Partial Differential Equations on Parameter-Dependent Domains Using Deep Neural Networks
 Abstract
-Partial differential equations (PDEs) are extensively utilized for modeling various physical phenomena.
-These equations often depend on certain parameters, necessitating either the identification of optimal parameters or solving the equation across multiple parameters to understand how a structure might react under different conditions.
+Partial differential equations (PDEs) are widely used for modeling various physical phenomena.
+These equations often depend on certain parameters, necessitating either the identification of optimal parameters or the solution of the equations over multiple parameters.
 Performing an exhaustive search over the parameter space requires solving the PDE multiple times, which is generally impractical.
-To address this challenge, Reduced Order Models (ROMs) are constructed from a snapshot dataset comprising parameter-solution pairs.
-ROMs facilitate the rapid solving of PDEs for new parameters.
-Recently, Deep Learning ROMs (DL-ROMs) have been introduced as a new method to obtain ROM.
-Additionally, the PDE of interest may depend on the domain, which can be characterized by parameters or measurements and may evolve with the system, requiring parametrization for ROM construction.
-In this paper, we develop a Deep-ROM capable of extracting and efficiently utilizing domain parametrization.
-Unlike traditional domain parametrization methods, our approach does not require user-defined control points and can effectively handle domains with varying numbers of components.
-Moreover, our model can derive meaningful parametrization even when a domain mesh is unavailable, a common scenario in biomedical applications.
-Our work leverages Deep Neural Networks to effectively reduce the dimensionality of the PDE and the domain characteristic function.
-Keywords Reduced Order Modeling  $\cdot$ Deep Learning  $\cdot$ Partial Differential Equations
+To address this challenge, reduced order models (ROMs) are built using a set of precomputed solutions (snapshots) corresponding to different parameter values.
+These models allow for a fast approximation of the PDE solution when new parameters are introduced.
+Recently, Deep Learning ROMs (DL-ROMs) have been introduced as a new method to obtain ROM, offering improved flexibility and performance.
+In many cases, the domain on which the PDE is defined also varies and may be described by parameters or measurements.
+Capturing this variation is important for building accurate ROMs but is often difficult, especially when the domain has a complex structure or changes topology.
+In this paper, we propose a Deep-ROM framework that can automatically extract useful domain parametrization and incorporate it into the model.
+Unlike traditional domain parameterization methods, our approach does not require user-defined control points and can effectively handle domains with varying numbers of components.
+It can also learn from domain data even when no mesh is available, which is often the case in biomedical applications.
+Using deep neural networks, our approach reduces the dimensionality of both the PDE solution and the domain representation, making it possible to approximate solutions efficiently across a wide range of domain shapes and parameter values.
+We demonstrate that our approach produces parametrizations that yield solution accuracy comparable to models using exact parameters.
+Importantly, our model remains stable under moderate geometric variations in the domain, such as boundary deformations and noise—scenarios where traditional ROMs often require remeshing or manual adjustment.
+Keywords Reduced Order Modeling $\cdot$ Deep Learning $\cdot$ Partial Differential Equations $\cdot$ Parameter-Dependent Domains
 1 Introduction
-Applications such as design optimization and real-world simulation often require solutions to partial differential equations across various parameters [1], [2], [3], [4].
-Typically, with given parameters $\bm{\lambda}$ the objective is to find a function $\bm{u}$ such that:
-$\mathcal{L}(\bm{u};\bm{\lambda})=f(\cdot;\bm{\lambda}),\text{ on $\Omega(\bm{\lambda})$}$ (1)
-$\mathcal{B}(\bm{u};\bm{\lambda})=g(\cdot;\bm{\lambda}),\text{ on $\partial\Omega(\bm{\lambda})$},$ (2)
-where $\mathcal{L}$ and $\mathcal{B}$ are differential operators, $f$ and $g$ are functions and $\Omega$ is the domain, and all of these can depend on the parameter vector $\bm{\lambda}$.
-This task can be formulated as finding a mapping from parameters to their corresponding solutions.
-A naive approach would be to just solve PDE for each parameter value we want to examine.
-This can often be prohibitive due to time and memory constraints.
-To mitigate this problem, a Reduced Order Model (ROM) is constructed, which relies on the assumption that all solutions for a specified parametric PDE belong to a lower-dimensional space.
-Classically, the problem of constructing ROM then involves determining the reduced basis [5].
-In order to construct the ROM, one needs to obtain a dataset of high-fidelity parameter-solution pairs.
-This is usually done by employing the Galerkin finite element method (FEM) [6] which approximates solutions within a finite-dimensional space and uses a weak formulation.
-To solve a single instance of a PDE, the solution is approximated in a finite-dimensional function space.
-The problem of solving the PDE is then reformulated as the problem of finding the coefficients of the basis functions in that space.
-This results in a large system of equations and the discretized solution of PDE can be written as a large vector.
-To solve for many parameter values, one constructs and uses ROM.
-Its development consists of two phases: offline phase in which the ROM is constructed, and online phase, where the solution for the new parameter is evaluated.
-To construct ROM, one collects the dataset $(\bm{\lambda},\bm{u}_{i}),i=1,\ldots,N$, where $\bm{\lambda}$ is the i-th vector of parameters and $\bm{u}_{i}$ is the full-order model (FOM) solution corresponding to these parameters.
-The dataset comes from measurements or high-fidelity numerical solvers.
-Usually, using this data, linear reduction techniques are used to project solutions and equations into lower-dimensional space to obtain the ROM.
-To get a solution for the new parameter, it is then only necessary to solve in lower-dimensional space, which is faster than solving the original PDE.
-However, evaluation still requires solving a smaller system of equations.
-Furthermore, classical ROMs depend on the quality of the linear projection.
-This may result in a high number of equations for a reduced model.
-Because of that, few nonlinear methods were developed [7], [8], [9].
-Regarding classical Reduced Order Models (ROMs), constructing ROMs that are effective across different domains presents several challenges.
-Firstly, to identify the projection onto the appropriate lower-dimensional space, all components of all solutions must have matching locations on the computational domain.
-This requirement is often unfeasible because irregular domains are discretized differently, making it impossible to ensure that all solutions have the same dimensionality.
-A common solution to this issue is to create a mapping between a fixed reference domain and the domain of interest, transforming the equation to apply to the fixed reference domain [10].
-However, this approach is highly problem-dependent and not universally applicable.
-Additionally, domains may be derived from measurements, have varying numbers of subdomains or holes, or lack an available mesh.
-In such cases, it is unclear how to parametrize these domains and construct an appropriate mapping to a reference domain.
-In this work, we address these challenges by employing convolutional autoencoder architecture on characteristic functions represented as a bitmap image to extract domain parametrization.
-This approach works in cases where the mesh is not available, when domains come from measurements, or when we have varying numbers of components or holes.
-The extracted parametrization is then incorporated into a Deep-ROM architecture.
-We demonstrate the effectiveness of our approach across various 2D problems.
+Applications such as design optimization and simulations in real-world settings often require solving partial differential equations (PDEs) for a wide range of parameter values [1], [2], [3], [4].
+Typically, given a parameter vector $\bm{\lambda}$, the goal is to find a solution $\bm{u}$ such that:
+$\mathcal{L}(\bm{u};\bm{\lambda})=f(\cdot;\bm{\lambda})\text{ on $\Omega(\bm{\lambda})$},$ (1)
+$\mathcal{B}(\bm{u};\bm{\lambda})=g(\cdot;\bm{\lambda})\text{ on $\partial\Omega(\bm{\lambda})$},$ (2)
+where $\mathcal{L}$ is a differential operator, $\mathcal{B}$ is a boundary operator, $f$ and $g$ are the source and boundary data, and $\Omega$ is a parameter-dependent domain.
+This task can be viewed as learning a solution map from the parameters $\bm{\lambda}$ to their corresponding solutions $\bm{u}$.
+A straightforward approach is to numerically solve the PDE for each parameter instance, but this becomes computationally prohibitive as the number of evaluations increases.
+To address this, reduced order models (ROMs) are commonly used to approximate the solution map with a much lower computational cost.
+These models are built on the assumption that solutions to the parametric PDE lie on a low-dimensional manifold [5].
+The construction of such a ROM typically involves two phases.
+In the offline phase, a dataset of high-fidelity parameter–solution pairs $(\bm{\lambda}_{i},\bm{u}_{i})$ is collected using methods such as the finite element method (FEM) [6] and ROM is obtained.
+In the online phase, the ROM is evaluated on unseen parameters.
+Full-order solutions are then projected on a reduced basis, often obtained using techniques such as Proper Orthogonal Decomposition (POD) [7].
+In the online phase, a reduced system is solved to obtain an approximate solution for a new parameter value, significantly reducing computation time.
+However, standard ROMs face limitations, especially when applied to PDEs defined on parameter-dependent domains.
+Constructing a projection-based ROM typically requires all solutions to be discretized over a common mesh or aligned computational domain, so that basis functions are defined consistently across parameter instances.
+This assumption breaks down when the domain itself varies with the parameters, which is common in many real-world scenarios.
+A typical workaround is to map each domain to a fixed reference configuration and transform the PDE accordingly [8].
+However, this mapping is often problem-specific and not easily generalized.
+Although this is feasible when domains vary smoothly and share the same topology, it becomes problematic when domains differ structurally, for instance, when holes appear or disappear, or when the number of connected components changes.
+In such cases, defining a single reference domain and constructing consistent mappings becomes difficult, or even impossible.
+To address these challenges, we propose a data-driven framework based on convolutional autoencoders to learn a compact domain representation directly from its characteristic function.
+We represent each domain as a binary image and use a convolutional autoencoder to extract a low-dimensional encoding.
+This encoding serves as a learned domain parametrization, which we then integrate into a neural reduced order model that predicts PDE solutions.
+This approach requires no meshing, no manual alignment, and no user-defined control points.
+It is naturally suited to problems where the domain is defined through measurements or where topological complexity varies across parameter instances.
+Moreover, in many real-world scenarios—such as biomedical simulations, flow over irregular surfaces, or domains derived from imaging—constructing a computational mesh is often impractical or infeasible.
+This limits the applicability of traditional ROM techniques that rely on consistent meshing, graph connectivity, or control-point mappings to handle domain variability.
+This challenge is particularly acute when domains are derived from imaging or experimental data, where mesh or graph structures may not be available at all.
+To overcome these limitations, we propose a fully data-driven, mesh-free approach that operates directly on binary images representing domain characteristic functions.
+Using convolutional autoencoders, our method learns compact geometric encodings without requiring explicit parametrization or meshing.
+We demonstrate the effectiveness of this approach on synthetic two-dimensional PDE problems designed to reflect practical challenges, including geometric variability, boundary noise, and structural differences across domains.
 2 Related work
@@
 2.1 Reduced Order Model via Proper orthogonal decomposition
-Proper orthogonal decomposition (POD) [15] is a linear reduction technique.
-The basic assumption is that even though discretized solutions belong to a high-dimensional vector space of dimension $m$, solutions can be well approximated by a lower-dimensional vector space of dimension $k\ll n$.
-In this setting, all solutions can be vectorized and assembled into a matrix $\mathbf{U}\in\mathbf{M^{mn}}$.
-To obtain the basis of such a lower-dimensional vector space, SVD decomposition [16] is calculated, which means that orthogonal matrices $\mathbf{\hat{U}}\in\mathbf{M^{m}},\mathbf{\hat{V}}\in\mathbf{M^{n}}$ orthogonal matrices and diagonal $\Sigma\in\mathbf{M^{mn}}$ with non-negative decreasing numbers on its diagonal $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{m}$ are found such that
+Proper orthogonal decomposition (POD) [7, 12] is a linear reduction technique.
+The basic assumption is that even though discretized solutions belong to a high-dimensional vector space of dimension $M$, solutions can be well approximated by a lower-dimensional vector space of dimension $k\ll M$.
+In this setting, all solutions can be vectorized and assembled into a matrix $\mathbf{U}\in\mathbf{\mathbb{R}^{M\times N}}$.
+To obtain the basis of such a lower-dimensional vector space, SVD decomposition [13] is calculated, which means that orthogonal matrices $\mathbf{\hat{U}}\in\mathbf{\mathbb{R}^{M\times M}},\mathbf{\hat{V}}\in\mathbf{\mathbb{R}^{N\times N}}$ and diagonal $\Sigma\in\mathbf{\mathbb{R}^{M\times N}}$ with non-negative diagonal elements $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{min\{M,N\}}$, are found such that
 $\mathbf{U}=\mathbf{\hat{U}\Sigma\hat{V}}$.
@@
 where $\bm{x}$ is $k$-dimensional vector.
-Now, instead of finding $\bm{u}$, we have to find $\bm{x}$.
-Then, when $Vx$ is substituted instead of $\bm{u}$ in the Galerkin FOM formulation, a new system with fewer unknowns is obtained.
+Now, instead of finding $\mathbf{u}$, we have to find $\mathbf{x}$.
+Then, when $\mathbf{Vx}$ is substituted instead of $\mathbf{u}$ in the Galerkin FEM formulation, a new system with fewer unknowns is obtained.
 In linear PDE, this results in decreasing the number of equations.
 In non-linear cases, this cannot be trivially done, so techniques such as the Discrete Empirical Interpolation Method (DEIM) [14] are developed.
-Further difficulties in employing this ROM arise when the domain depends on parameters because all discretized solutions need to have the same dimensionality and each coordinates of all solutions have to correspond to the same points in the domain.
-To obtain ROM on parametric domains $\Omega(\bm{\lambda})$, equations are mapped on reference domain $\hat{\Omega}$.
+Further difficulties in employing this ROM arise when the domain depends on parameters because all discretized solutions need to have the same dimension and each coordinates of all solutions have to correspond to the same points in the domain.
+To obtain ROM on parametric domains $\Omega(\bm{\lambda})$, equations are mapped to the reference domain $\hat{\Omega}$.
 Mapping $\mathcal{M}(\cdot;\bm{\lambda})$ maps the reference domain into $\Omega(\bm{\lambda})$, so the correspondence of points across different domains is established.
@@
 Control points have to be chosen carefully so that they capture the shape characteristics well.
-This imposes difficulties when one wants to obtain one ROM for situations where there are a varying number of components or holes.
+This imposes difficulties when one wants to obtain a single ROM for situations where there are a varying number of components or holes.
 2.2 Physical Informed Neural Networks for Reduced Order Model
@@
 2.3 Deep ROM (DL-ROM)
-N.
 Franco et al.
@@
 The model consists of three neural networks: an encoder $\Psi_{E}^{S}$, a decoder $\Psi_{D}^{S}$, and a multi-layer perceptron (MLP) [28] $\Phi_{S}$.
-To obtain ROM, the encoder-decoder pair is trained to minimize reconstruction error:
+To obtain ROM, the encoder-decoder pair (i.e.
+autoencoder) is trained to minimize reconstruction error:
 $\frac{1}{N}\sum_{i=1}^{N}\|\bm{u}_{i}-\Psi_{D}^{S}(\Psi_{E}^{S}(\bm{u}_{i}))\|^{2}.$
-The encoder for standard DL-ROM $\Psi_{E}^{S}(\bm{u}_{i})$ maps the solution to a lower-dimensional space, called encoding, and the decoder lifts encoding to the original solution size.
-The autoencoder is then trained to minimize reconstruction error.
-Finally, $\Phi_{S}$ is trained to minimize the mean-squared error between its output and encoder encoding:
+The encoder for standard DL-ROM $\Psi_{E}^{S}(\bm{u}_{i})$ maps the solution to a lower-dimensional space, called encoding, and the decoder lifts the encoding to the original solution size.
+Finally, $\Phi_{S}$ is trained to minimize the mean square error between its output and encoder encoding:
 $\frac{1}{N}\sum_{i=1}^{N}\|\Phi_{S}(\bm{\lambda}_{i})-\Psi_{E}^{S}(\bm{u}_{i})\|^{2}.$
-Solution $\bm{u}$ for a new value of parameter $\bm{\lambda}$ is then:
+The solution $\bm{u}$ for a new value of parameter $\bm{\lambda}$ is then:
 $\bm{u}=\Psi_{D}^{S}(\Phi_{S}(\bm{\lambda})).$ (3)
-The whole architecture is shown on Figure 1
+The entire architecture is shown on Figure 1
 Figure 1: Basic DL-ROM architecture: encoder $\Psi_{E}^{S}$ and decoder $\Psi_{D}^{S}$ are trained to minimize reconstruction error between full order model solution and its reconstruction, After training, encodings of solutions are calculated with encoder $\Psi_{E}^{S}$ and Multilayer perceptron (MLP) $\Phi_{S}$ is trained to map parameters to encoding.
-To obtain a solution for new parameters $\bm{\lambda}$, parameters are passed through $\Phi_{S}$ and then through the decoder $\Psi_{D}^{S}$.
-Notice that getting solutions from this kind of reduced-order model doesn’t require solving any additional equations and can be efficiently done for multiple parameters at the same time, suing GPU parallel computing.
-In this work, Convolutional Neural Networks (CNN) [32] were used for the encoder and decoder and all domains involved were rectangular.
+To obtain a solution for a new parameter $\bm{\lambda}$, parameters are passed through $\Phi_{S}$ and then through the decoder $\Psi_{D}^{S}$.
+Notice that getting solutions from this kind of reduced-order model does not require solving any additional equations and can be efficiently done for multiple parameters at the same time, using GPU parallel computing.
+Authors used Convolutional Neural Networks (CNN) [29] for the encoder and decoder and all domains involved were rectangular.
 2.4 Deep Learning Proper Orthogonal Decomposition ROM
@@
 Columns of matrix $\mathbf{A}$ are $\bm{a}_{i},i=1,2,\ldots N$.
-The essence of the POD method is to get orthonormal vectors $\bm{\Phi_{i}},i=1,2,\ldots k$, $k\ll N$ vectors that maximize:
-$\frac{1}{N}\sum_{i=1,\ldots N,j=1,\ldots k}|\langle\bm{a}_{i},\bm{\Phi}_{j}\ \rangle|^{2}$
-So, the following approximation hold:
+The essence of the POD method is to get orthonormal vectors $\bm{\Phi_{i}},i=1,2,\ldots k$, $k\ll N$ that maximize:
+$\frac{1}{N}\sum_{i=1,\ldots N,j=1,\ldots k}|\langle\bm{a}_{i},\bm{\Phi}_{j}\ \rangle|^{2}.$
+Then, the following approximation hold:
 $\bm{u}\approx\mathbf{V}_{k}\bm{u}_{k}$
 Matrix $\mathbf{V}_{k}$ is usually found using SVD decomposition or randomized SVD (rSVD) [31].
-So, the solution $u$ can be approximated with POD basis and represented as $k$-dimensional vector $[\langle\bm{u},\bm{\Phi}_{1}\rangle,\ldots\langle\bm{u},\bm{\Phi}_{k}\rangle]$.
+The solution $u$ can be approximated with POD basis and represented as $k$-dimensional vector $\langle\bm{u},\bm{\Phi}_{1}\rangle,\ldots\langle\bm{u},\bm{\Phi}_{k}\rangle]$.
 Then, similar to the black-box DL-ROM, the model consists of three networks: encoder $\Psi_{E}$, decoder $\Psi_{D}$, and parameter-encoding network $\Phi$.
@@
 Using the proposed autoencoder technique, one can obtain ROM for non-rectangle domains using the GNN autoencoder.
-In the proposed approach, graph pooling and a fully-connected layer work only when all domains have the same number of nodes and domain parametrization is known as well.
+However, graph pooling and a fully-connected layer work only when all domains have the same number of nodes and domain parametrization is known as well.
 In applications, especially in biomedicine, both constraints may not hold.
+In [37], a harmonic mapping is used to map the domain and solution to a reference configuration.
+The neural operator is learned on a reference configuration and mapped to the original domain via inverse of the harmonic mapping.
+This approach may not be feasible when such harmonic mapping is difficult or impossible to construct due to a wide geometric or topological variability.
+Similarly as in our work, in [38] a characteristic function of the domain is used, and the solutions are interpolated onto a rectangular grid using CNN.
+The input is a domain characteristic function and the output is the solution on that domain.
+Contrary to our work, this model is not capable of handling non-geometric parameters simultaneously.
 3 Our method
 In this work, we develop an efficient Reduced Order Model (ROM) that operates effectively across different domains, based on a Deep Learning Reduced Order Model (DL-ROM) approach.
-We refer to our method as Deep Learning domain aware ROM (DL-DA ROM).
-To achieve this, meaningful parametrization of domains is essential.
-Domains can be described parametrically, may evolve as part of the system (e.g., moving domains), or can be obtained from measurements.
-Since the ROM for given parameters has to predict the domain and values of solutions on that domain, we separate those two tasks.
-In this work, we assume that the solution is at least continuous and can be interpolated onto rectangular mesh without significant loss of accuracy, see Figure 2.
-By interpolating in this manner, Convolutional neural networks (CNNs) can be utilized to construct an autoencoder for solutions.
-This approach is particularly beneficial in scenarios where the domain is derived from measurements and meshing is prohibitively expensive.
-Figure 2: Sample interpolated onto equidistant rectangular grid (on the left). Sample (on the right).
+We refer to our method as Deep Learning Domain Aware ROM (DL-DA-ROM).
+A key part of this approach is learning a meaningful parametrization of the domain.
+In many problems, the domain may be defined through parameters, evolve with the system (e.g., moving domains), or come from measurements.
+Since the ROM needs to predict both the domain and the solution on that domain, we separate those two tasks.
+We assume that the solution is at least continuous and can be interpolated onto rectangular grid without significant loss of accuracy, see Figure 2.
+This interpolation allows us to use CNNs to build an autoencoder for the solutions.
+This setup is especially useful when domains come from measurement data and creating a mesh is either difficult or too costly.
+While it is possible to interpolate the solution and then mask out the region outside the domain, we found that using the full interpolated solution yields better performance.
+Interpolation naturally smooths the data, reducing sharp discontinuities introduced by irregular meshes or complex geometries.
+This smoothing effect makes the solution easier to approximate with CNNs, which are particularly well-suited for capturing smooth spatial patterns.
+As a result, CNN-based autoencoders can more efficiently learn a compact representation of the solution field without being disrupted by artificial discontinuities at domain boundaries.
+Figure 2: Left panel shows a sample function, and the right panel shows its interpolation into an equidistant rectangular grid.
 When domains are parametrically described, it is not necessary to assume that each domain has the same number of parameters; they may have varying numbers of subdomains or holes.
-Our approach assumes that each domain characteristic function can be interpolated onto a rectangular grid, see Figure 3.
-Figure 3: Mesh representation of the domain (left), an image of domain characteristic function (right).
-The domain of the problem can be derived either from measurements or parameters.
-In our proposed method, its characteristic function is described as a bitmap image, with pixels assigned a value of one or zero depending on whether the pixel is inside or outside the domain.
-A convolutional autoencoder is trained using binary cross-entropy loss [40] to classify each pixel as part of the domain or not which is given by
+Our approach assumes that each domain characteristic function can be interpolated onto a fixed rectangular grid, see Figure 3.
+Figure 3: Comparison between a mesh-based domain representation (left) and the corresponding interpolated characteristic function represented as an image used by our method (right).
+The binary image representation allows us to learn a low-dimensional domain encoding without requiring a mesh.
+To represent the domain, we use its characteristic function interpolated onto a rectangular grid, resulting in a binary image.
+This image is stored as a bitmap, where each pixel is assigned a value of 1 if the corresponding point lies inside the domain and 0 otherwise.
+A convolutional autoencoder is then trained on these bitmap images to extract a compact latent representation, which serves as a learned, low-dimensional domain parameterization.
+This mesh-free approach accommodates complex geometries and topological variations without requiring control points, meshing, or predefined mappings.
+To extract the domain parametrization, we train a convolutional autoencoder on these binary images, using the binary cross-entropy loss[39] to classify each pixel as part of the domain or not which is given by
 $l(x,y)=\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}-y_{ij}\log(x_{ij})-(1-y_{ij})\log(1-x_{ij}),$
@@
 The encoded domains are then used as the parametrization of the domain.
-The domain autoencoder is inspired by semantic segmentation encoder-decoder models in computer vision [41], [42], with two classes indicating whether a pixel is inside or outside the domain.
+The low-dimensional latent code produced by the encoder is used as domain parametrization in our model.
+The autoencoder architecture is inspired by the encoder-decoder networks commonly used in semantic segmentation [40, 41], with two classes that distinguish whether each pixel lies within or outside the domain.
 This approach can be extended to problems with multiple domains, such as fluid-structure interaction problems, providing simultaneous geometric parametrization for the entire system.
-If domains can be generated synthetically, it is possible to use more samples for the domain autoencoder than for the solution autoencoder.
-Domain autoencoder is shown in Figure 4.
+If domains are synthetically generated, it is also possible to train the domain autoencoder with a larger dataset than the solution autoencoder, which can help improve the quality of the learned parametrization and its generalization to unseen geometries.
+The domain autoencoder is shown in Figure 4.
 Figure 4: Domain Autoencoder used for domain parametrization.
-Input to the convolutional neural network is a characteristic function represented as a bitmap image.
-The output of the autoencoder is a rectangular grid with the probability that a given pixel belongs to the domain.
-The encoder output is used as domain parametrization.
-If the domain is either known from measurements or is parametrized, the encoding of the domain can be used as additional parameters in the network $\Phi$.
-The model is shown in Figure 5 and all important terms in this work are given in table 1.
-[sca]
-Figure 5: Deep ROM for parametric domains: In addition to vanilla Deep ROM, we introduce domain autoencoder, which extracts parameters for the domain.
-These encodings are then used as additional parameters in Deep ROM.
-label meaning $\Omega(\bm{\lambda})$ domain of the problem, in our work, is dependent on parameters $\Psi_{E}^{S}$ solution encoder $\Psi_{D}^{S}$ solution decoder $\Phi_{S}$ neural network that maps parameters into solution encoding $\Psi_{E}^{D}$ domain encoder $\Psi_{D}^{D}$ domain decoder
+The input to the convolutional neural network is a characteristic function represented as a bitmap image.
+The output of the autoencoder is a rectangular grid where each pixel value represents the predicted probability that the pixel belongs to the domain.
+The low-dimensional encoder output (latent representation) is used as the domain parametrization.
+Fully-connected neural network $\Phi_{S}$ is trained to map input parameters (e.g., domain encoding and physical parameters) to the solution representation.
+The predicted latent solution is then passed through the decoder $\Psi^{S}_{D}$ to reconstruct the full solution on the interpolated grid.
+The model is shown in Figure 5 and all important terms are given in Table 1.
+Figure 5: Deep ROM architecture for PDEs with parameter-dependent domains.
+In addition to the standard Deep ROM structure, our method incorporates a domain autoencoder (bottom) that encodes the characteristic function of the domain into a low-dimensional latent representation.
+This representation is combined with other problem parameters and passed to a fully connected network $\Phi_{S}$, which predicts the solution encoding.
+The predicted latent solution is then decoded using $\Psi^{S}_{D}$ to reconstruct the solution on the full-order model (FOM) grid.
+Both the solution and domain autoencoders are trained using CNNs.
+The domain representation serves as a learned geometric parametrization that allows the ROM to generalize across different domain shapes and configurations.
+label meaning $\Omega(\lambda)$ parameter-dependent domain $\Psi_{E}^{S}$ solution encoder $\Psi_{D}^{S}$ solution decoder $\Phi_{S}$ neural network that maps parameters into solution encoding $\Psi_{E}^{D}$ domain encoder $\Psi_{D}^{D}$ domain decoder
 Table 1: Notation used in this work.
-Constructing the ROM, also known as the offline phase, involves gathering the data, either with numerical simulation, measurements, or possibly a combination of both.
-Domain and solution autoencoders are trained to obtain meaningful low-dimensional encodings.
+Constructing the ROM-also known as the offline phase, involves gathering the data, either using numerical simulation, measurements, or possibly a combination of both.
+During this phase, both the domain and solution autoencoders are trained to obtain meaningful low-dimensional representations.
 Before training the solution autoencoder training solution data is standardized by applying the following transformation:
-$\bm{u}_{i}=\frac{\bm{u}_{i}-\mu_{s}}{\sigma_{s}},1,\ldots N$
-where $N$ is a number of samples in the training dataset, $\mu_{s}$ and $\sigma_{s}$ are mean and standard deviation of the training dataset After training the solution and domain autoencoders, domain encoding, and other equation parameters are concatenated.
-Training parameters are standardized by applying the following transformation:
-$\lambda_{ij}=\frac{\lambda{ij}-\mu_{pj}}{\sigma_{pj}},i=1,\ldots N,j=1\ldots k$
-where $k$ is dimension of parameter vector, $N$ number of samples and $\mu_{pj},\sigma_{pj}$ are mean and standard deviation of each parameter separately.
-Afterward, the network $\Phi_{S}$ is trained to predict solution encodings from parameters.
-The procedure is summarized in Algorithm 1.
-After all data collection and training are completed, solutions for unseen parameters are calculated in the online phase, as detailed in Algorithm 2.
-For new values of equation parameters and domain, this involves passing the domain through the domain encoder $\Psi_{E}^{D}$ concatenating the resulting encoding with the equation parameters, passing them through the MLP $\Phi_{S}$ and finally through the solution decoder $\Psi_{D}^{S}$ to obtain prediction $\hat{u}$.
-The final ROM solution is then given by:
-$\bm{u}=\sigma_{s}\bm{u}+\mu_{s}$
+$\bm{u}_{i}=\frac{\bm{u}_{i}-\mu_{s}}{\sigma_{s}},1,\ldots N,$
+where $N$ is the number of samples in the training dataset and $\mu_{s}$ and $\sigma_{s}$ are the mean and standard deviation of the training dataset.
+After the solution and the domain autoencoders are trained, we evaluate the domain encoder to extract the low-dimensional domain parameters.
+These are then concatenated with the PDE parameters.
+Prior to training the solution prediction network $\Phi_{S}$, the concatenated input equation and domain parameters are standardized:
+$f_{ij}=\frac{f_{ij}-\mu_{j}^{p}}{\sigma_{j}^{p}},i=1,\ldots N,j=1\ldots k,$
+where $f_{i,j}$ the $j$-th component of the concatenated parameter vector for the $i$-th training sample, $k$ is dimension of a parameter vector, $N$ number of samples and $\mu_{j}^{p},\sigma_{j}^{p}$ are mean and standard deviation of the $j$-th component of parameter vectors in the training set.
+The network $\Phi_{S}$ is then trained to map from this combined, standardized parameter vector to the corresponding solution encoding.
+The full procedure is summarized in Algorithm 1.
+Once the training is complete, solutions for unseen parameters are calculated in the online phase, as detailed in Algorithm 2.
+Given a new domain and PDE parameters, the domain is first passed through the trained encoder $\Psi_{E}^{D}$, the standardized result is concatenated with the standardized PDE parameters, and the solution is reconstructed using the trained networks.
+The predicted ROM solution $\hat{u}$ is then transformed back to the original scale using:
+$\bm{\hat{u}}=\sigma_{s}\bm{\hat{u}}+\mu_{s}.$
 Algorithm 1 Offline phase
@@
 Save all trained networks
-Save mean $\mu_{s}$ and standard deviation $\sigma_{s}$ of training dataset and mean $\mu_{p}$ and standard deviation $\sigma_{p}$ of training feature dataset
+Save mean $\bm{\mu}_{s}$ and standard deviation $\bm{\sigma_{s}}$ of training dataset and mean $\bm{\mu}_{p}$ and standard deviation $\bm{\sigma}_{p}$ vectors of training feature dataset
 Algorithm 2 Online phase
-: Neural networks $\Psi_{D}^{S},\Psi_{E}^{D}$, and $\Phi_{S}$, values of parameters and/or domains for which we want to evaluate solution $(\bm{\lambda}_{i},\bm{c}_{i}),i=1,\ldots M$, mean and standard deviation of training parameters $\bm{f}_{i}$ $\mu_{p},\sigma_{p}$, mean and standard deviation of solutions $\mu_{s},\sigma_{s}$
+: Neural networks $\Psi_{D}^{S},\Psi_{E}^{D}$, and $\Phi_{S}$, values of parameters and/or domains $\bm{d}$ for which we want to evaluate solution $(\bm{\lambda}_{i},\bm{c}_{i}),i=1,\ldots M$, mean and standard deviation of training parameters $\bm{f}_{i}$ $\bm{\mu}_{p},\bm{\sigma}_{p}$, mean and standard deviation of solutions $\mu_{s},\sigma_{s}$
 Solutions for given parameters and domains $\bm{u}_{i},i=1,\ldots M$
 Compute domain parameters $\bm{d}_{i}=\Psi_{E}^{D}(\bm{c}_{i})$
-Concatenate $f_{i}=(\bm{\lambda}_{i},\bm{d}_{i}),i=1,\ldots M$
+Concatenate $\bm{f}_{i}=(\bm{\lambda}_{i},\bm{d}_{i}),i=1,\ldots M$
 Standardize $\bm{f}_{i}$ based on training mean and standard deviation
 Calculate ROM prediction $\bm{\hat{u}}=\Psi_{D}^{S}(\Phi_{S}(\bm{f}_{i}))$
-Return solution $\sigma_{s}\bm{\hat{u}}+\mu_{s}$
-Experiments and Results
-Through this section, we will use relative error to evaluate our approach. The relative error $\epsilon$ for sample $(\bm{\lambda},\bm{u}_{h},\bm{d}_{h})$ is:
-$\epsilon(\bm{u}_{h},\hat{\bm{u}_{h}})=\frac{\sqrt{\sum_{i,j}(u^{i,j}_{h}d^{i,j}_{h}-\hat{u}^{i,j}_{h}d^{i,j}_{h})^{2}}}{\sqrt{\sum_{i,j}(u^{i,j}_{h})^{2}d^{i,j}_{h}}},$
-where $\bm{\lambda}$ is the parameter vector, $\bm{u}_{h}$ solution discretized on a mesh with $n\times m$ points, $\bm{\hat{u_{h}}}$ is DL-ROM prediction and $\bm{d}_{h}$ is an approximation of characteristic functions of the domains on that same mesh.
-Relative error at i-th row and j-th column is:
-$\epsilon_{i,j}(u_{k},\hat{u}_{h})=\frac{u_{h}^{i,j}d_{h}^{i,j}-\hat{u}^{i,j}_{h}d^{i,j}_{h}}{\sqrt{\sum_{k,l}(u^{k,l}_{h})^{2}d^{k,l}_{h}}},$
+Return solution $(\sigma_{s}\bm{\hat{u}}+\mu_{s})\odot d$
+4 Experiments and Results
+Throughout this section, we will use relative error to evaluate our approach. The relative error $\epsilon$ for a sample $(\bm{\lambda},\bm{u}_{h},\bm{d}_{h})$ is:
+$\epsilon(\bm{u}_{h},\hat{\bm{u}_{h}})=\frac{\sqrt{\sum_{i,j}(u^{i,j}_{h}-\hat{u}^{i,j}_{h})^{2}d^{i,j}_{h}}}{\sqrt{\sum_{i,j}(u^{i,j}_{h})^{2}d^{i,j}_{h}}},$
+where $\bm{\lambda}$ is the parameter vector, $\bm{u}_{h}$ solution discretized on a mesh with $n\times m$ points, $\bm{\hat{u_{h}}}$ is DL-ROM prediction and $\bm{d}_{h}$ is an approximation of characteristic functions of domains on that same mesh.
+For visualization purposes, we also define a pointwise relative error at position $(i,j)$:
+$\epsilon_{i,j}(u_{k},\hat{u}_{h})=\frac{(u_{h}^{i,j}-\hat{u}^{i,j}_{h})d^{i,j}_{h}}{\sqrt{\sum_{k,l}(u^{k,l}_{h})^{2}d^{k,l}_{h}}},$
 where $u^{i,j}_{h},\hat{u}^{i,j}_{h},d^{i,j}_{h}$ are discretized solutions, model approximation of the solution, and characteristic function value in i-th row and j-th column.
-We use FEM implementation in FreeFEM [43] to obtain solutions and Pytorch [44] for defining and training neural networks.
-All solutions are projected on a rectangular equidistant mesh, in order to use CNN.
-In all examples, Leaky ReLU is used as an activation function after convolution.
-We use batch normalization [45] after each layer, except for the last layers of the encoder and decoder.
-In all examples, the number of channels per layer in the decoder mirrors the encoder architecture.
-Upsampling operations are used only in some layers, such that the resolution of outputs in the decoder layers mirrors the resolution of outputs in the encoder layers.
-For all networks, Adam optimizer [46] with One Cycle Policy scheduling [47] is used that adjusts the learning rate at each training step, with a maximum of 1e-3 for the encoder-decoder network.
+We use a FreeFEM++ software [42] to generate finite element solutions and PyTorch [43] to implement and train all neural networks.
+All solutions are projected onto a uniform rectangular grid to enable the use of CNNs.
+For the advection–diffusion examples, the elliptical-hole and circular-hole cases are discretized on a $128\times 128$ grid, while the Navier–Stokes case uses a $256\times 512$ grid.
+Interpolation relative error is calculated as L2 relative error of the data on interpolated mesh projected to original mesh and original data and is approximately $0.06\%$ for every example.
+In all experiments, we use Leaky ReLU activations in convolutional layers and apply batch normalization [44] after each layer, except for the final layers of the encoder and decoder.
+The decoder mirrors the encoder in architecture and number of channels.
+Upsampling layers are used in selected layers to ensure output resolutions match those of the corresponding encoder layers.
+All networks are trained using Adam optimizer [45] with the One Cycle Policy learning rate scheduler [46], with a maximum learning rate of $1\times 10^{-3}$ for autoencoder networks.
 All training, test, and validation sets used have $1000$ samples.
 The batch size for the autoencoder is $50$.
-Autoencoder network hyperparameters were chosen by trial and error with respect to autoencoder reconstruction, where we searched on latent dimensions $15,20,30$ and the number of channels for each encoder layer (the decoder mirrors the encoder).
-Tried hyperparameters are shown in Table 2.
-Autoencoders with more channels per layer had better training and test accuracy, so we picked the best two for further analysis.
-Latent dimension $[15,20,30]$
-Nbr. of kernels in encoder $[4,4,8,8,16,16,16]$, $[4,8,16,32,32,64,64]$ , $[8,16,16,32,64,64,64]$
-Learning rate 1e-3
+Autoencoder network hyperparameters are chosen by trial and error with respect to autoencoder reconstruction, where we searched on latent dimensions $15$, and $30$ and the number of convolutional channels for each encoder layer (the decoder mirrors the encoder).
+The hyperparameters tested are shown in Table 2, are identical across all examples.
+Autoencoders with more channels per layer have better training and test accuracy, so we picked the best two for further analysis.
+Latent dimension $15,30$
+Nbr. of channels in each
+encoder layer
+$[4,4,8,8,16,16,16]$,
+$[4,8,16,32,32,64,64]$ ,
+$[8,16,16,32,64,64,64]$
+Strides
+$[1,1,2,2,2,1,1]$
+Maximum learning rate 1e-3
 Batch size 50
@@
 Table 2: Autoencoder grid search hyperparameters.
-We trained 9 models with varying channel sizes and latent dimension sizes.
+We trained 6 models with varying channel sizes and latent dimension sizes.
 The number of channels in decoder layers mirrors the encoder number of channels per layer.
-We have chosen two autoencoders with the smallest reconstruction errors on the test set, both with $8,16,16,32,64,64,64,64$ channels in encoders and dimensions $20$ and $30$ for further analysis.
-We have not encountered any overfitting problems by increasing autoencoder capacity.
-However, neural network $\Phi_{S}$ easily overfits.
-Since $\Phi_{S}$ is usually small compared to autoencoder and trains fast, we employed grid search to choose the best hyperparameters for $\Phi_{S}$.
+We have chosen two autoencoders with the smallest mean relative reconstruction errors on the validation set, both with $8,16,16,32,64,64,64$ channels in encoders and dimensions $15$ and $30$ for further analysis.
+Since $\Phi_{S}$ is usually small compared to autoencoder and fast to train, we employ grid search to choose the best hyperparameters for $\Phi_{S}$.
 Hyperparameters are shown in Table 3.
 The grid search was done for two ROM representations, obtained from solution autoencoders with $8,16,16,32,64,64,64$ channels per layer and latent dimensions equal to $15$ and $30$, respectively.
-Domain features are obtained from autoencoders trained on $50000$ domains with $32,32,64,64,64,64,64,64$ channels per layer in each layer and strides $1,1,2,2,2,1,1,1$ per layer respectively.
+Domain features are obtained from autoencoders trained on $9000$ domains with $8,16,16,32,64,64,64$ channels per layer in each layer and strides $1,1,2,2,2,1,1$ per layer, respectively.
 For each model, the One Cycle Policy scheduler with Adam optimizer was used.
-We have trained two different domain autoencoders, one with latent dimension $15$ and one with latent dimension $30$.
-We choose the best hyperparameters with respect to the mean-squared error between the prediction of $\Phi_{S}$ and ROM representation.
+We trained two different domain autoencoders, one with latent dimension $20$ and one with latent dimension $30$.
+We choose the best hyperparameters with respect to the mean-squared error between the prediction of $\Phi_{S}$ and ROM representation on validation set.
 Hyperparameter Values
@@
 Number of layers 1,2,3,4
-Neurons per layer 64, 128, 256, 512, 1204, 2048
+Neurons per layer 64, 128, 256, 512, 1024, 2048
 Dropout rate 0, 0.025, 0.05, 0.1, 0.2
@@
 Batch size $8,16,32,64,128$
-Table 3: Hyperparameters grid search for $\Phi$ is performed, which results in total $4200$ models.
+Table 3: Hyperparameters grid search for $\Phi$ is performed, which results in $4200$ models in total.
 4.1 Advection-diffusion on different domains
-Advection-diffusion equation on parametric domains:
-$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ without elipse}\\ u=1&\text{ on edge of ellipse}\\ u=0&\text{for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (4)
+In this section, we evaluate the accuracy and generalization capabilities of our proposed DL-DA-ROM method on a parameterized advection-diffusion problem.
+The goal is to compare the performance of models using exact domain parameters (e.g., ellipse descriptors) versus those using domain encodings learned by the autoencoder.
+This setting allows for a controlled comparison, as both the physical and geometric parameters are known.
+By analyzing prediction accuracy across different latent dimensions, we assess the quality of the learned domain representations and their ability to replace traditional parameterizations.
+Precisely, we consider the following advection-diffusion equation on parametric domains:
+[…]
 Parameters are:
   • transport angle $\phi\in[0,2\pi]$ which determines vector $\mathbf{c}=(cos(\phi),sin(\phi))$
-  • center of ellipse $(x_{0},y_{0})\in[0.1,0.9]^{2}$
-  • angle of major axis of ellipse and x-axis $\alpha\in[0,\pi]$
-  • $\beta\in[1,10]$
-  • half-axes for ellipse are then taken randomly from interval depending on the center and angle of the ellipse
-In total, there are $7$ parameters.
-In this example, we know exact geometric parametrization, so we can obtain DL-ROM performance with calculated and exact parametrization and directly compare those two.
-For each domain parametrization, we have trained a fully-connected neural network $\Phi_{S}$ with hyperparameters determined by grid search.
-Comparison of vanilla DL-ROM and our approach is shown in Table 4.
+  • center of the ellipse $(x_{0},y_{0})\in[0.15,0.85]^{2}$
+  • angle between major axis of the ellipse and the x-axis $\alpha\in[0,\pi]$
+  • boundary flux $\beta\in[1,10]$
+  • half-axes $a,b\in[0.05,0.25]$
+In total, there are $7$ parameters: two PDE parameters—transport angle $\phi$ and boundary flux $\beta$—and five geometric parameters defining the ellipse: center coordinates $(x_{0},y_{0})$, half-axes $a$ and $b$, and orientation angle $\alpha$.
+We assume that the ellipse has to be at least 0.01 units away from each of the boundaries to avoid geometric artifacts.
+Since the domain is generated from known ellipse parameters, we can directly compare the performance of our model using either exact geometric parameters or those inferred by the domain autoencoder.
+For each type of domain parametrization, we train a fully connected neural network $\Phi_{S}$ with hyperparameters selected via grid search.
+A comparison of the vanilla DL-ROM and our approach is shown in Table 4.
 We compare DL-ROM performance when exact ellipse parameters are used with domain parametrization calculated with the autoencoder.
-In all cases, models with calculated ellipse parameters perform better on the test set than models with exact ellipse parameters.
-In fact, we report that more than $1300$ models out of $4200$ trained in grid search, performed better on parameters of domain obtained autoencoder than the best model trained on exact domain parameters, for each combination.
-In this case, it can be seen that vanilla DL-ROM is more sensitive to ROM dimension, than our approach.
-Results are shown on Figure 6.
-test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact ellipse parameters only 0.0554552 0.0554552 0.0605261 0.0605261 exact and calculated elipse parameters 0.0454250 0.0425800 0.0458577 0.0438216 calculated elipse parameters 0.0465742 0.0452926 0.0451946 0.0449190
-Table 4: Comparison of DL-ROM and DL-DA ROM performance: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension.
-We differentiate between cases where geometric parameters are exactly known (1st row), when we add domain-encoded features to exact geometric features (2nd row), and when we use only geometric parameters obtained by autoencoder (3rd row).
-We can see that using autoencoder domain parametrization is beneficial for model performance.
+Performance of the best models is shown in Figure 6, evaluated on one of the examples from the test set.
+test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact ellipse parameters only 0.0348 0.0348 0.0383 0.0383 calculated elipse parameters 0.0339 0.0361 0.0310 0.03975
+Table 4: Comparison of DL-ROM and DL-DA ROM performance: Difference of the performance on relative test errors depending on the ROM encoding dimension and domain encoding dimension.
+We differentiate between cases where geometric parameters are exactly known (1st row) and when we use only geometric parameters obtained by autoencoder (2nd row).
+We can see that using the autoencoder domain parametrization does not degrade model performance.
 We report on the best model for each ROM encoding dimension and domain encoding dimension.
-Models trained on calculated domain parameters perform better than models trained using only exact domain parameters.
-Figure 6: Top row: Exact solution, left, network prediction for a model trained on exact domain parameters (center), network prediction for a model trained on approximated domain parameters.
-Bottom row: relative error for a model trained on exact domain parameters (left), relative error for a model trained on approximated domain parameters (middle), relative error for a model trained on both exact and calculated domain parameters (right).
-Results are shown for ellipse with center $(0.76,0.38)$, half-axes $a=0.14,b=0.11$ and angle $\alpha=1.51$, $\beta=6.92$, $\mu=0.29$.
+Models trained on calculated domain parameters perform similar as models trained using only exact domain parameters.
+Even though results are similar, larger fully-connected neural network is needed in case of calculated domain parameters (4 layers with 2048 neurons vs 4 layers with 512 neurons).
+Figure 6: Comparison of predictions and relative errors for the DL-ROM with ROM encoding dimension $30$ and domain encoding dimension $20$.
+From left to right: Exact solution, network prediction for a model trained on exact domain parameters, network prediction for a model trained on approximated domain parameters, relative error for the model trained on exact domain parameters, and relative error for the model trained on approximated domain parameters.
+Results correspond to the ellipse with center (0.815, 0.456)(0.815,\,0.456), half-axes $a=0.194$, $b=0.110$, transport angle $\phi=2.497$, and boundary flux $\beta=6.884$.
 All models were trained on $1000$ samples.
+In order to compare the exact and autoencoder domain parametrization, we investigate error distributions for models with the best performance according to Table 4.
+Histograms of relative errors are in Figure 7.
+The results indicate that domain encoding improves model robustness, as reflected by a reduced presence of outliers.
+Figure 7: Histograms of relative errors on test set for model with ROM encoding dimension 30, trained on the exact elipse parameters (blue) and model with ROM encoding dimension 30, trained on 20-dimensional domain encoding (orange).
+When using exact parametrization, $15$ samples have the relative error larger than $0.1$ and $213$ larger than $0.05$.
+In case of using calculated domain parametrization, $15$ samples have the relative error larger than $0.1$ and $121$ larger than $0.05$.
+There are more error outliers when using exact ellipse parametrization.
 4.2 Domains with varying number of parameters
-We experiment again on the advection-diffusion equation on parametric domains.
-In this case, domains can depend on a different number of parameters.
-Such a domain may be difficult to parametrize by classical techniques.
-We are investigating domains that are squares with $1$ to $4$ circular holes.
-Equations are:
-$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ whitout circles}\\ u=2&\text{ on edge of circles}\\ u=0&\text{ for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (5)
+We experiment again with the advection-diffusion equation on parametric domains.
+In this setting, the domains vary in complexity by depending on a different number of parameters, which makes them challenging to parametrize using classical techniques.
+Specifically, we consider square domains that contain between one and four circular holes.
+More precisely, we consider the following boundary value problem:
+[…]
 Parameters are:
@@
   • $m\in[1,4]$ number of circular holes
-  • centers of holes $(x_{i},y_{i})$, $i=1,\ldots m$.
-  • circles radii $r_{i}$, $i=1,\ldots m$.
-  • $\beta\in[1,10]$
+  • centers of holes $(x_{i},y_{i})$, $i=1,\ldots m$
+  • circles’ radii $r_{i}$, $i=1,\ldots m$
+  • boundary flux $\beta\in[1,10]$
+In total, we again have two PDE parameters: transport angle $\phi$ and boundary flux $\beta$, but there are between three and twelve geometric parameters, namely, the radius and center coordinates for each circular hole.
 To generate domains, first, we randomly choose the number of circles $m$.
-Then we randomly choose the centers of circles so the distance between the centers and the square boundary is at least $0.1$.
+Then we randomly choose the centers of the circles so that the distance between the centers and the outer boundary is at least $0.1$.
 Finally, radii are chosen randomly so that the distance between all circles is at least $0.1$ and $r\in[0.05,0.2]$.
 This procedure ensures that circles do not overlap and that all circles are of reasonable size.
-We have chosen the same encoder-decoder architectures as in the previous example and after training solution autoencoders, a grid search was performed to obtain $\Phi_{S}$ hyperparameters.
-Results are shown in Table 5 and Figure 7.
-A larger ROM dimension mostly does not affect DL-ROM performance, while a smaller domain parametrization encoding dimension is beneficial in this case.
-test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 calculated domain parameters 0.0239904 0.0264217 0.0209590 0.0267886
-Table 5: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension.
+We have chosen the same encoder-decoder architectures as in the previous example, and after training solution autoencoders, a grid search was performed to obtain $\Phi_{S}$ hyperparameters.
+Table 5 shows the average relative error on the test set for different combinations of domain and solution latent dimensions, while Figure 8 shows an example of a solution corresponding to a geometric configuration from the test set.
+test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 calculated domain parameters 0.0252 0.0242 0.0244 0.0234
+Table 5: Test relative errors obtain usign different ROM and the domain encoding dimensions.
 In this case, domains have varying numbers of parameters.
@@
 We report on the best model for each ROM encoding dimension and domain encoding dimension.
-Figure 7: Exact solutions (left), model prediction (middle), relative error (right) for model trained on $1000$ random samples for parameters $\beta=0.977$, $\mu=0.842$, with centers $(0.23,0.73),(0.57,0.22),(0.25,0.37),(0.79,0.79)$ and radii $0.11,0.10,0.14,0.10$ respectively.
-Furthermore, we test how sensitive DL-DA ROM DL-ROM is on small perturbations of the circle boundary.
-The sensitivity of this model is especially important if we train on synthetically generated data and test only on possibly imprecisely measured domains and parameters.
-Instead of testing the trained model on a domain with circular holes, we test the model on predicting the correct solution when the domain is a unit square without ellipses.
-The idea is that these ellipses model slightly deformed circles.
-The correct solution is again generated using the finite element method.
-We tested on solutions with parameters $\beta=1.19617,\theta=1.71657$ with $4$ holes, with centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively.
-A comparison between FEM and our DL-ROM is shown in Figure 8.
-We tested how sensitive our model is to perturbations of the boundary.
-In these cases, instead of circular holes, we have elliptic holes.
-The relative error for increasing deformation is given in Table 6.
-Ratio between small and large half-axis two half-axes in % $\epsilon$
-1 0.0101
-0.95 0.0120
-0.9 0.0142
-0.8 0.0166
-Table 6: Relative errors for the same $\beta$, $\theta$ and ellipse centers as in 8 ($\beta=1.19617$, $\theta=1.71657$).
-Ellipses have centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$, radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively and varying rations between small and large half-axis.
-We can see an increase in relative error with higher deformation of the domain boundary, but the increase is not rapid.
-Figure 8: Exact solutions (left) not multiplied by characteristic function, model prediction (middle) not multiplied by characteristic function, relative error (right) multiplied by characteristic function for slightly deformed circles with parameters $\beta=1.19617$, $\theta=1.71657$.
-Ellipses have centers, $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively.
-Small half-axes are $20\%$ smaller than large half-axes for each ellipse.
-Even though the model was not trained on elliptic holes, we can see a small relative error further from the boundary.
-The error is the largest at the boundary since the solution autoencoder did not capture ellipses in the original training dataset.
-This demonstrates that the model can work with small perturbations of the boundary.
-Conclusion
-In this work, we present the extension of DL-ROM to various 2D domains.
-This approach can be extended to 3D problems.
-Our algorithm can be used in problems where domains come from measurements.
-However, this comes at the expense of the main disadvantage of this approach, which is the first step of interpolating into a rectangular domain.
-In some problems, such as 3D thin pipes with torsion, rectangular data can be much larger than the actual number of nodes required to describe the pipe, which can be a computational bottleneck.
-In future work, we plan to mitigate this issue as well.
-We demonstrated that meaningful geometrical parametrization can be obtained using neural networks, which may be useful in other reduced-order model techniques as well.
-Acknowledgments
-This research was performed using the advanced computing service provided by the University of Zagreb University Computing Centre - SRCE.
+A smaller domain encoding dimension is beneficial in this case.
+Figure 8: Exact solutions (left), model prediction (middle), relative error (right) for model trained on $1000$ random samples for parameters $\beta=1.671$, $\phi=0.182$, with centers $(0.322,0.379)$ and $(0.374,0.736)$, and radii $0.130$ and $0.124$, respectively.
+Robustness to Domain Deformations.
+To test the robustness of the model, we evaluated DL-ROMs trained on domains with circular holes using test domains in which the circles were slightly deformed into ellipses.
+Specifically, each circle was transformed by reducing its minor semiaxis to (1p)(1-p) times the major semiaxis, where $p$ is a deformation factor.
+To introduce variability, even-numbered ellipses were aligned along the $x$-axis, and odd-numbered ones were aligned along the $y$-axis.
+This setup simulates realistic domain perturbations, such as geometric noise or measurement uncertainty.
+The trained models were tested on these deformed domains, and their predictions were compared with finite element solutions.
+Table 6 reports the average test errors for increasing levels of deformation.
+As expected, the error increases as the deformation grows.
+However, the model does not break down and continues to produce reasonable approximations even when tested on domains with types of geometries not seen during training.
+This suggests that the method generalizes well to moderate domain perturbations.
+Figure 9 shows one such out-of-distribution example.
+The relative error is highest near the boundaries of the ellipses, where the geometry most significantly deviates from the training distribution.
+Still, the model captures the global solution behavior accurately.
+The histogram of test errors in Figure 10 confirms that even under moderate distribution shifts, no significant outliers appear, indicating that the predictions remain stable and that the method is robust.
+test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 $p=0$ 0.0252 0.0242 0.0244 0.0234 $p=0.05$ 0.0260 0.0252 0.0252 0.0254 $p=0.1$ 0.0280 0.0275 0.0278 0.0294 $p=0.15$ 0.0309 0.0310 0.0319 0.0356
+Table 6: Relative test errors for DL-ROMs trained on domains with circular holes ($p=0$), evaluated on domains with increasingly deformed elliptical holes.
+Here, 1p1-p denotes the ratio between the smaller and larger semi-axis of the ellipse.
+We report results for the best model for each ROM and domain encoding dimension.
+Performance degrades gradually with increasing deformation, but remains acceptable, demonstrating robustness to moderate boundary changes.
+Figure 9: The exact solution (left) not multiplied by characteristic function, the model prediction (middle) not multiplied by a characteristic function, relative error (right) multiplied by characteristic function for slightly deformed circles with parameters $\beta=9.643$, $\phi=4.129$.
+Ellipses have centers, $(0.601,0.0481)$, $(0.441,0.0.206)$, $(0.0.789,0.0.206)$, and $(0.334,0.770)$, and radii of non-deformed circles $0.113$, $0.105$, $0.101$, and $0.118$, respectively.
+Small half-axes are $15\%$ smaller than large half-axes for each ellipse.
+While relative error is highest near domain boundaries due to unseen shapes during training, the model maintains low error elsewhere, showing robustness to domain perturbation.
+Figure 10: Histogram of relative errors on test set for model with ROM encoding dimension 30 and domain encoding dimension 30. Only $38$ out of $1000$ samples have relative error larger than $0.05$.
+4.3 Navier-Stokes equations
+In this section, we evaluate the DL-DA-ROM framework on the Navier–Stokes equations, a significantly more complex and challenging problem.
+The Navier–Stokes equations govern fluid flow and include significant nonlinearities, providing a challenging and realistic scenario to test the effectiveness and robustness of our methodology in practical applications.
+The considered scenario involves fluid flow through a cavity with a geometrically complex boundary, motivated by the simulations of blood flow through vessels with aneurysms.
+Specifically, the domain is characterized as a channel with a variable width, representing an aneurysm-like enlargement, and further complicated by the presence of a circular obstacle to introduce additional geometric complexity.
+We vary the Reynolds number between 1 and 100.
+Furthermore, to assess the robustness of our model, we deliberately test parameter configurations that lie outside the original training distribution.
+This allows us to verify that our model can handle moderate extrapolation scenario, ensuring stability and reasonable accuracy for previously unseen, but not drastically different, geometries and flow conditions.
+Precisely, we consider the following boundary value problem:
+[…]
+where the domain $\Omega(\lambda))$ is a cavity without a circle.
+The length of the domain is $L=4.5$ and $\sigma$ is the fluid stress tensor.
+Top side is deformed as $R+he^{-w(x-1.25)^{2}}+s(x)$, where the spline $s$ is given by $9$ points $(0,0),(0.25,0),(x_{s1},y_{s1}),(x_{s2},0),(2.2,0),(x,he^{(-w(-L/2)^{2})}-he^{-w(x-L/2)^{2}})$ for $x=3.0,3.5,4.0,4.5$ and the bottom is deformed as $he^{w(x-1.25)^{2}}-s(x)$.
+The parameters are:
+  • undeformed pipe width $R\in[0.1,0.2]$
+  • obstacle radius $rCirc\in[0.1,0.15]$
+  • $x_{s1}\in[0.5,1]$
+  • $y_{s1}\in[0.15,0.25]$
+  • $x_{s2}\in[1.5,2]$
+  • obstacle center $(x_{0},y_{0})\in[0.5,1.5]\times[0.05,0.1]$
+  • inflow amplitude $a\in[5,15]$
+  • $w\in[0.05,0.2]$
+  • $h\in[0.15,0.2]$
+  • $Re\in[1,100]$
+In total, we have two PDE parameters: inflow amplitude $a$ and Reynolds number $Re$ and $9$ geometric parameters: pipe width $R$, obstacle radius $rCirc$, obstacle center $(x_{0},y_{0})$, spline points $x_{s1},y_{s1},x_{s2}$ and $w$ and $h$ determining the shape of the cavity.
+The solutions are interpolated on the $256\times 512$ grid on $0.5,0.725]\times[0,4.5]$.
+We train the model to predict the velocity in $x$ and $y$-direction.
+The autoencoder is trained on the velocity components in both the $x$ and $y$-directions, while all reported relative errors correspond to the velocity magnitude.
+A comparison of model performance using calculated versus exact domain parameters is presented in Table 7, with the representative results shown in Figure 11.
+We emphasize that despite the increased complexity of both the governing equations and the domain geometry compared to the advection-diffusion case, the average error only slightly increases and remains well within acceptable limits for reduced-order modeling.
+test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact domain parameters only 0.0412 0.0412 0.0335 0.0335 calculated domain parameters 0.0361 0.0348 0.0379 0.0333
+Table 7: Test relative errors obtained using different ROM and the domain encoding dimensions.
+We differentiate between cases where geometric parameters are exactly known (1st row) and when we use only geometric parameters obtained by the autoencoder (2nd row).
+We can see that the use of the autoencoder domain parameterization does not negatively affect DL-ROM performance.
+We report on the best model for each ROM encoding dimension and domain encoding dimension.
+Most Models trained on calculated domain parameters perform better than the models trained using only exact domain parameters.
+Even though results are similar, larger fully-connected neural network is needed in case of calculated domain parameters (4 layers with 2048 neurons vs 4 layers with 512 neurons).
+Figure 11: The exact solutions (top left), model prediction (top right) and relative error (bottom) for model trained on $1000$ random parameter samples, with 2020-dimensional domain parametrization obtain using $a=10.34$ and $Re=30.1$.
+We further evaluate the sensitivity of the model to domain perturbations by testing parameter configurations that fall outside the range of the original training set.
+Specifically, we consider domains with wider cavities ($w\in[0.2,0.25]$), using the model with ROM and domain encoding dimensions both set to $30$.
+Additionally, we increase the Reynolds number slightly beyond the training range, with $Re\in[100,110]$, while other parameters remain within the original bounds.
+A total of $100$ such samples were generated, resulting in an average relative error of approximately $0.065$.
+We closely examine outlier cases and find that one source of increased error occurs when the domain autoencoder fails to accurately reconstruct the domain geometry (see Figure 12).
+Nevertheless, even in such cases, the method provides a reasonable approximation with a relative error of $0.138$.
+Conversely, Figure 13 illustrates a case where the domain encoding performed well, resulting in a relative error of $0.087$.
+These examples indicate that, despite some reconstruction issues, the model remains robust and produces acceptable solutions even under moderate parameter and geometry shifts.
+Figure 12: Out of distribution sample with the highest relative error of $0.138$.
+Domain reconstruction (top left), correct solution (top right), prediction (bottom left), relative error (bottom right).
+Domain autoencoder is not able to properly reconstruct the domain, producing artificial holes in the domain.
+This implies that calculated domain parametrization is not appropriate for this domain.
+However, the model still finds a reasonable solution.
+Figure 13: Typical out of distribution sample.
+Domain reconstruction (top left), correct solution (top right), prediction (bottom left), relative error (bottom right).
+In this case, the domain autoencoder is able to reconstruct the domain, thus able to extract more meaningful representation.
+The distribution of the relative error on the test set for the best model reported in Table 7 is shown in Figure 14. We can see that the model predictions are stable across the test dataset.
+Figure 14: Histogram of relative errors on test set for the model with ROM encoding dimension 30 trained with exact domain parameters (blue), and trained with domain encoding dimension 30 (orange).
+5 Conclusion
+We presented a method for building reduced-order models for PDEs on a parameter-dependent domains, including scenarios where the domain geometry is not explicitly given but instead defined through images or measurements.
+Unlike many existing ROM techniques that require consistent meshing or graph-based representations, our approach directly utilizes bitmap images of domain characteristic functions.
+This makes it particularly suitable for practical applications where mesh generation is costly, unreliable, or unavailable—such as simulations derived from imaging data or experimental measurements.
+Our results demonstrate that the learned domain encoding is not only compact but also meaningful; in cases with known geometric parameters, our encoding achieved similar accuracy to models using exact parameterizations.
+Furthermore, our method effectively handled significant variations in domain shape, changes in the number of features (such as holes), and moderate boundary perturbations without losing stability.
+Although the use of domain parametrization obtained via the CNN autoencoder successfully enables mapping from physical parameters to the latent solution space, we observed that this configuration generally requires a larger network capacity to achieve comparable accuracy.
+This behavior likely stems from a partial misalignment between the domain parametrization and the reduced-order latent space, which increases the effective nonlinearity of the mapping.
+Future work will focus on improving the consistency between the autoencoder latent representation and the parameter space, for example through joint or coupled training strategies or latent regularization terms.
+Such approaches could reduce the required network size and improve generalization while preserving the interpretability and robustness of the proposed DA-DL-ROM framework.
+Conceptually, our approach extends naturally to three-dimensional problems.
+However, from a practical perspective, computational efficiency remains an important challenge due to the increased size of the input grids.
+Improving computational efficiency for three-dimensional implementations is currently a focus of our ongoing work.
+Additionally, the learned domain representations may offer utility beyond DL-ROM frameworks, potentially benefiting other situations requiring domain parameterization where explicit parameters are unavailable or in different machine learning-based model reduction contexts.
+6 Acknowledgments
+This research was carried out using the advanced computing service provided by the University of Zagreb University Computing Centre - SRCE.
 This research was supported by the Croatian Science Foundation under the project number IP-2022-10-2962 (I.M., B.M.
 and D.V.).
+M.B.
+has been partly supported by the National Science Foundation via grants NSF DMS-2208219 and NSF DMS-2205695.